# RobotKeyframing: Learning Locomotion with High-Level Objectives via Mixture of Dense and Sparse Rewards

### v1

@@ Abstract
Keywords: Legged Robots, Multi-Critic Reinforcement Learning, Motion Imitation
@@ 1 Introduction
Legged robots hold a great promise for becoming household companions [1] or automated performers in the entertainment industry [2, 3]. In these applications, it is crucial for robot controllers to perform natural and directable behavior from simple high-level user command inputs beyond the common commands used in the robotic domain such as joystick velocity commands [4, 5] or target base position [6, 7].
@@
Inspired by character animation, we aim to equip legged robots with more refined control by incorporating sparse and temporal high-level objectives as keyframes. The primary goal of this work is to develop a locomotion controller that enables the robot to fulfill specified partial or full-pose targets while infilling natural behavior during the intermediate periods. This goal aligns with recent advancements in using reinforcement learning (RL) for legged robots due to their promising robustness and flexibility [10, 11]. However, learning a policy that accurately meets keyframes without imposing undesired constraints at intermediate periods presents challenges, particularly due to the need to handle sparsity in the keyframe objectives. Acquiring effective policies requires a meticulous reward design procedure that carefully balances these sparse rewards with other dense rewards which are crucial for regularizing and encouraging natural motion.
In this work, we present a novel framework that unifies timed high-level objectives with natural locomotion of legged robots through temporal keyframes. Along with the imitation objective similar to Peng et al. [12] for natural motion generation, our pipeline allows specifying full or partial high-level targets, including base position, orientation, and joint postures. We propose using a multi-critic RL framework to address the challenge of managing groups of sparse and dense rewards by learning distinct value functions. Our method also employs a novel transformer-based architecture to encode a variable number of goals with arbitrary time intervals. Unlike typical sequence-to-sequence transformers [13], we propose a lightweight sequence-to-token module that can be used autoregressively within a feedback control loop. We demonstrate the effectiveness of our framework through experiments both in simulation and on real-world hardware. Our policies successfully guides the robot to meet multiple keyframes at the required times, for both position and posture targets. Furthermore, the multi-critic approach showcases better convergence with less hyperparameter tuning compared to the conventional single-critic method. Our experiments also reveals that using a transformer-based encoder to anticipate future goals significantly enhances goal-reaching accuracy.
The contribution of this paper is threefold: (i) We introduce RobotKeyframing, a novel learning-based framework for integrating high-level objectives in natural locomotion of legged robots; (ii) We propose using multi-critic RL to handle the mixture of dense and sparse rewards and a novel sequence-to-token encoder to accommodate a variable number of keyframes; (iii) We validate the effectiveness of our method through extensive experiments in simulation and on hardware.
@@ 3.1 Problem Setup
To integrate high-level control objectives into the robotic control framework, we employ sparse keyframes that require a robot to achieve specific goals at predetermined times. Each keyframe contains a full or partial combination of a variety of targets such as global base position $\hat{\bm{p}}\in\mathbb{R}^{3}$, global base orientation $(\hat{\phi},\hat{\zeta},\hat{\psi})\in\mathbb{R}^{3}$ where $\phi$, $\zeta$, $\psi$ denote roll, pitch, and yaw angles respectively, and full posture specified by joint angles $\hat{\bm{\theta}}_{j}\in\mathbb{R}^{N_{j}}$ where $N_{j}$ is the number of joints. Each keyframe is also assigned with a specific time $\hat{t}\in\mathbb{R}$ in the future at which the robot is expected to meet the goals. In summary, the high-level objectives are specified through these keyframes $\bm{K}=\left(\bm{k}^{1},\bm{k}^{2},...,\bm{k}^{n_{k}}\right)$, where $\bm{k}^{i}=\left(\hat{\bm{g}},\hat{t}\right)^{i}$ and $\hat{\bm{g}}\subset\left\{\hat{\bm{p}},\hat{\phi},\hat{\zeta},\hat{\psi},\hat{\bm{\theta}}_{j}\right\}$. Here, $n_{k}\leq N_{k}$ where $n_{k}$ and $N_{k}$ denote the actual and the maximal number of keyframes, respectively. We aim to support an arbitrary number of keyframes, allowing for the flexible specification of high-level objectives only as needed. The main goal is to train a locomotion policy for legged robots that not only meets these keyframes but also maintains a natural style in the intervals between them. To avoid undesired restrictions on the intermediate periods, policy’s task performance is evaluated exclusively at the designated times, making the keyframe objectives temporally sparse. However, relying solely on keyframes to train the control policy may result in undesirable motions. Thus, it is crucial to have additional rewards for regularizing and promoting a natural motion style. In this regard, we incorporate AMP [12] as a general style guide for the robot, encouraging the policy to behave naturally and similarly to an offline motion dataset from real animals [44]. The style and regularization rewards are evaluated at every step of the episode, making them temporally dense. The mixture of sparse and dense rewards presents a unique challenge that is difficult to manage effectively with standard RL frameworks. Further details on the observation, action, reward definitions, and training procedure can be found in Appendix A.
@@ 3.2 Multi-Critic RL for Dense-Sparse Reward Mixture
Modern RL algorithms [45, 46, 47] typically employ the actor-critic paradigm, where the actor decides the action to take, and the critic evaluates the action by estimating the value function. To effectively manage a complex mixture of temporally dense and sparse rewards, we employ a multi-critic (MuC) RL framework by Martinez-Piazuelo et al. [48] as shown in Figure 2. It involves training a set of critic networks $\{V_{\phi_{i}}\}_{i=0}^{n}$ to learn distinct value functions associated with different reward groups $\{r_{i}\}_{i=0}^{n}$. Similar concepts have been used to balance a set of dense rewards [49, 50]; however, we aim to adapt the multi-critic method to the context of dense and sparse reward combination. We design each reward group to contain either exclusively dense or sparse rewards. This division is essential for effectively managing the distinct temporal characteristics of each reward type and facilitates value estimation.
@@ 3.3 Transformer-based Keyframe Encoding
In our system, each input token corresponds to a particular keyframe. At every time step $t$, each keyframe $\bm{k}^{i}$ is transformed spatially and temporally into a robot-centric view, resulting in a goal error $\Delta\bm{g}^{i}_{t}$ and a calculated time to goal $\hat{t}^{i}-t$. These are then concatenated with the robot state $\bm{s}_{t}$ to form a single token. Additionally, we incorporate a self-goal keyframe, $\bm{x}^{0}_{t}$, as the first token in the sequence. This token represents a state with zero error and zero time to goal, which ensures that the control system remains operational despite the absence of active goals or after achieving all goals. The transformer encoder receives the sequence of tokens $\bm{X}_{t}=(\bm{x}^{0}_{t},...,\bm{x}^{n_{k}}_{t})$, where $\bm{x}^{0}_{t}=(\bm{s}_{t},\bm{0},0)$, and $\bm{x}^{i}_{t}=(\bm{s}_{t},\Delta\bm{g}^{i}_{t},\hat{t}^{i}-t)$ for $i=1,...,n_{k}$.
In scenarios where the number of active keyframes is less than the maximum capacity of the system, we apply masking to ignore the surplus tokens and focus only on the relevant keyframes. Furthermore, we also apply masking to keyframes once their designated time is reached and surpassed by a few steps. This practice prevents past goals from inappropriately influencing the long-term behavior of the policy. The output from the transformer encoder is then forwarded to a max-pooling layer, which condenses the encoded goal features for delivery to the subsequent multilayer perceptrons (MLP). By leveraging transformer’s ability to handle sequences of varying lengths, our architecture can effectively integrate multiple and arbitrary numbers of goals into the control process.
@@ 4 Results
4.1 Keyframe Tracking
We demonstrate that our trained policy effectively reaches keyframes at the designated times through several simulation experiments. Given keyframes consisting of position goals, our policy reaches its targets with notable precision, as illustrated in Figure 4(a) by the horizontal trajectories for two example scenarios with different number of keyframes. Furthermore, our framework offers control over target reaching time and can generate diverse behaviors for the same targets by specifying different time profiles. This is depicted in Figure 4(b) through snapshots of robot motion when provided with keyframes consisting of the same position goal, but different target times. Full posture targets are also supported along with position and orientation goals. Figure 5 shows snapshots of the robot motion given different keyframe scenarios, highlighting that our policy accurately meets its full posture targets and maintains a natural style while reaching them.
@@ 5 Discussion
Conclusion: This paper presents RobotKeyframing, a learning-based control framework designed to incorporate high-level objectives into the natural locomotion of legged robots through a sequence of keyframes. Simulation and hardware experiments demonstrate the efficacy of our framework. The sparse reward imposed by keyframe objectives is effectively handled by a multi-critic PPO algorithm. In addition, the transformer-based architecture is adaptive to an arbitrary number of target keyframes and improves accuracy in reaching targets through future goal anticipation.
@@ A.1 Observation and Action Space
The observation of the policy is composed of two main components: state observation and goal observation. State observation at time $t$ include the linear velocity ($\bm{v}$) and angular velocity ($\bm{\omega}$) of the base in local coordinates, current joint angles ($\bm{\theta}_{j}$), current joint velocities ($\bm{\dot{\theta}}_{j}$), projected gravity in the base frame ($\bm{g}_{proj}$), base height ($h$) and previous actions ($\bm{a}_{prev})$,
@@ A.2 Reward Terms
Joint soft limits $\mathcal{K}\left(\max\left(\bm{\theta}_{j}-\bm{\theta}_{j,min},\bm{\theta}_{j,max}-\bm{\theta}\right),0.1,0\right)$
@@
Goal position $\Phi^{i}\left(\mathcal{K}(\bm{p}-\bm{\hat{p}}^{i},0.2,0)\right)$
Goal roll $\Phi^{i}\left(\mathcal{K}(\phi-\hat{\phi}^{i},0.1,0)\right)$
Goal pitch $\Phi^{i}\left(\mathcal{K}(\zeta-\hat{\zeta}^{i},0.1,0)\right)$
Goal yaw $\Phi^{i}\left(\mathcal{K}(\psi-\hat{\psi}^{i},0.3,0)\right)$
Goal posture $\Phi^{i}\left(\mathcal{K}(\left\lVert\bm{\theta}_{j}-\hat{\bm{\theta}}_{j}^{i}\right\rVert,0.2,0)\right)$
@@ A.3 Dataset Preparation
We use a database of motion capture from dogs introduced by Zhang et al. [44]. The motions are retargeted to the robot skeleton using inverse kinematics for the end-effectors’ positions with some local offsets to compensate for the different proportions of the robot and dog. A subset of around 20 minutes of data was used, removing the undesired motions such as smelling the ground, walking on slopes, etc. We augment this dataset with other motion clips animated by artists to include more diversity in the dataset. The frame rate is adjusted to that of the simulation, i.e. 50 frames per second.
@@ A.4 Training Procedure
We utilize Isaac Gym [55] for simulating the physical environment. At the start of each episode, the robot is either set to a default state or initialized according to a posture and height sampled from the dataset with Reference State Initialization (RSI). RSI plays a crucial role in capturing and learning the specific style of motion, as highlighted in previous studies such as Peng et al. [56]. Keyframes are derived either randomly or directly from a reference data trajectory. Our methodology incorporates a learning curriculum, beginning with keyframes entirely sourced from reference data and progressively increasing the proportion of randomly generated keyframes. To generate random keyframes, we start by selecting a time interval for each goal within a predetermined range. Subsequently, the distance and direction of the target position relative to the previous goal (or the initial position for the first goal) are sampled based on a specified range. The yaw angle is also chosen from a set range and adjusted relative to the previous goal. The robot’s full posture is sampled from the dataset to ensure the target posture is feasible. The roll, pitch, and height of the keyframe are aligned with the corresponding attributes of the target posture frame.
The meticulous sampling of target keyframes is critical for ensuring their feasibility and preventing them from impeding effective policy learning.
@@ A.5 Hardware Implementation Details
Domain randomization is added during training to achieve a robust policy that can be executed on hardware. Similar to Kang et al. [58], we randomize friction coefficients, motor stiffness and damping gains and actuator latency. Furthermore, we add external pushes during training. Although joint limits are softly taken into account in the simulation, we found it crucial to terminate episodes when reaching joint limits to ensure a stable deployment on hardware. We use a motion capture system to receive the global position and orientation of the robot. These are used to compute the relative errors to the target goals and are then passed to the policy. Other observations are computed based on the outputs from the state estimator.
A.6 Future goal anticipation

### v2

@@ Abstract
Project website: https://sites.google.com/view/robot-keyframing
Keywords: Legged Robots, Multi-Critic Reinforcement Learning, Transformer, Motion Imitation
@@ 1 Introduction
Legged robots hold a great promise for becoming household companions [1] or automated performers in the entertainment industry [2, 3]. In these applications, it is crucial for robot controllers to perform natural and directable behavior from simple high-level user command inputs beyond the typical commands used in the robotic domain such as joystick velocity commands [4, 5] or target base position [6, 7].
@@
With inspiration from the character animation technique, we aim to equip legged robots with more refined control by incorporating sparse and temporal high-level objectives as keyframes. The primary goal of this work is to develop a locomotion controller that enables the robot to fulfill specified partial or full-pose targets while infilling natural behavior during the intermediate periods. This goal aligns with recent advancements in using reinforcement learning (RL) for legged robots due to their promising robustness and flexibility [10, 11]. However, learning a policy that accurately meets keyframes without imposing undesired constraints at intermediate periods presents challenges, particularly due to the need to handle sparsity in the keyframe objectives. Acquiring effective policies requires a meticulous reward design procedure that carefully balances these sparse rewards with other dense rewards which are crucial for regularizing and encouraging natural motion.
In this work, we present a novel framework that unifies timed high-level objectives with natural locomotion of legged robots through temporal keyframes. Along with the imitation objective similar to Peng et al. [12] for natural motion generation, our pipeline allows specifying full or partial high-level targets, including base position, orientation, and joint postures. We propose using a multi-critic RL framework to address the challenge of managing groups of sparse and dense rewards by learning distinct value functions. Our method also employs a novel transformer-based architecture to encode a variable number of goals with arbitrary time intervals. Unlike typical sequence-to-sequence transformers [13], we propose a lightweight sequence-to-token module that can be used autoregressively within a feedback control loop. We demonstrate the effectiveness of our framework through experiments both in simulation and on real-world hardware. Our policies successfully guide the robot to meet multiple keyframes at the required times, for both position and posture targets. Furthermore, the multi-critic approach showcases better convergence with less hyperparameter tuning compared to the conventional single-critic method. Our experiments also reveals that using a transformer-based encoder to anticipate future goals significantly enhances goal-reaching accuracy.
The contribution of this paper is threefold: (i) We introduce RobotKeyframing, a novel learning-based framework for integrating high-level objectives in natural locomotion of legged robots; (ii) We propose using multi-critic RL to effectively handle the mixture of dense and sparse rewards, along with a sequence-to-token encoder to accommodate a variable number of keyframes; (iii) We validate the effectiveness of our method through extensive experiments in simulation and on hardware.
@@ 3.1 Problem Setup
To integrate high-level control objectives into the robotic control framework, we employ sparse keyframes that require a robot to achieve specific goals at predetermined times. Each keyframe contains a full or partial combination of a variety of targets such as global base position $\hat{\bm{p}}\in\mathbb{R}^{3}$, global base orientation $(\hat{\phi},\hat{\zeta},\hat{\psi})\in\mathbb{R}^{3}$ where $\phi$, $\zeta$, $\psi$ denote roll, pitch, and yaw angles respectively, and full posture specified by joint angles $\hat{\bm{\theta}}_{j}\in\mathbb{R}^{N_{j}}$ where $N_{j}$ is the number of joints. Each keyframe is also assigned with a specific time $\hat{t}\in\mathbb{R}$ in the future at which the robot is expected to meet the goals. In summary, the high-level objectives are specified through these keyframes $\bm{K}=\left(\bm{k}^{1},\bm{k}^{2},...,\bm{k}^{n_{k}}\right)$, where $\bm{k}^{i}=\left(\hat{\bm{g}},\hat{t}\right)^{i}$ and $\hat{\bm{g}}\subset\left\{\hat{\bm{p}},\hat{\phi},\hat{\zeta},\hat{\psi},\hat{\bm{\theta}}_{j}\right\}$. Here, $n_{k}\leq N_{k}$ where $n_{k}$ and $N_{k}$ denote the actual and the maximal number of keyframes, respectively. We aim to support an arbitrary number of keyframes, allowing for the flexible specification of high-level objectives only as needed.
The main goal is to train a locomotion policy for legged robots that not only meets these keyframes but also maintains a natural style during the intervals between them. To avoid undesired restrictions on the intermediate periods, policy’s task performance is evaluated exclusively at the designated times, making the keyframe objectives temporally sparse. However, relying solely on keyframes to train the control policy may result in undesirable motions. Thus, it is crucial to have additional rewards for regularizing and promoting a natural motion style. In this regard, we incorporate AMP [12] as a general style guide for the robot, encouraging the policy to behave naturally and similarly to an offline motion dataset from dogs [44]. The style and regularization rewards are evaluated at every step of the episode, making them temporally dense. The mixture of sparse and dense rewards presents a unique challenge that is difficult to manage effectively with standard RL frameworks. Details on the observation, action, reward definitions, and training setup can be found in Appendix A.
@@ 3.2 Multi-Critic RL for Dense-Sparse Reward Mixture
Modern RL algorithms [45, 46, 47] typically employ the actor-critic paradigm, where the actor decides the action to take, and the critic evaluates the action by estimating the value function. To effectively manage a complex mixture of temporally dense and sparse rewards, we employ a multi-critic (MuC) RL framework by Martinez-Piazuelo et al. [48] as shown in Figure 2. It involves training a set of critic networks $\{V_{\phi_{i}}\}_{i=0}^{n}$ to learn distinct value functions associated with different reward groups $\{r_{i}\}_{i=0}^{n}$. Similar concepts have been used to balance a set of dense rewards [49, 50]; however, we extend the multi-critic method to the context of dense and sparse reward combination, where we demonstrate its distinct advantages. We design each reward group to contain either exclusively dense or sparse rewards. This division is essential for effectively managing the distinct temporal characteristics of each reward type and facilitates value estimation.
@@ 3.3 Transformer-based Keyframe Encoding
In our system, each input token $\bm{x}^{i}_{t}$ is a vector corresponding to a particular keyframe. At every time step $t$, each keyframe $\bm{k}^{i}$ is transformed spatially and temporally into a robot-centric view, resulting in a goal error $\Delta\bm{g}^{i}_{t}$ and a calculated time to goal $\hat{t}^{i}-t$. These are then concatenated with the robot state $\bm{s}_{t}$ to form a single token. Additionally, we incorporate a self-goal keyframe, $\bm{x}^{0}_{t}$, as the first token in the sequence. This token represents a state with zero error and zero time to goal, which ensures that the control system remains operational despite the absence of active goals or after achieving all goals. The transformer encoder receives the sequence of tokens as a matrix $\bm{X}_{t}=[\bm{x}^{0}_{t};\dots;\bm{x}^{n_{k}}_{t}]$, where $\bm{x}^{0}_{t}=[\bm{s}_{t},\bm{0},0]$, and $\bm{x}^{i}_{t}=[\bm{s}_{t},\Delta\bm{g}^{i}_{t},\hat{t}^{i}-t]$ for $i=1,\dots,n_{k}$.
In scenarios where the number of active keyframes is less than the maximum capacity of the system, we apply masking to ignore the surplus tokens and focus only on the relevant keyframes. Furthermore, we also apply masking to keyframes once their designated time is reached and surpassed by a few steps. This practice prevents past goals from inappropriately influencing the long-term behavior of the policy. The output from the transformer encoder is then forwarded to a max-pooling layer, which condenses the encoded goal features for delivery to the subsequent multilayer perceptrons (MLP). By leveraging transformer’s ability to handle sequences of varying lengths, our architecture can effectively integrate multiple and variable numbers of goals into the control process.
@@ 4 Results
4.1 Keyframe Matching
We demonstrate that our trained policy effectively reaches keyframes at the designated times through several simulation experiments. Given keyframes consisting of position goals, our policy reaches its targets with notable precision, as illustrated in Figure 4(a) by the horizontal trajectories for two example scenarios with different number of keyframes. Furthermore, our framework offers control over target reaching time and can generate diverse behaviors for the same targets by specifying different time profiles. This is depicted in Figure 4(b) through snapshots of robot motion when provided with keyframes consisting of the same position goal, but different target times. Full posture targets are also supported along with position and orientation goals. Figure 5 shows snapshots of the robot motion given different keyframe scenarios, highlighting that our policy accurately meets its full posture targets while maintaining a natural style. Further quantitative evaluation of keyframe matching is provided in Section A.6.
@@ 5 Discussion
Conclusion: This paper presents RobotKeyframing, a learning-based control framework designed to incorporate high-level objectives into the natural locomotion of legged robots through a sequence of keyframes, leveraging transformer models and multi-critic reinforcement learning. Through simulation and hardware experiments, we demonstrated the effectiveness of our approach: The sparse reward imposed by keyframe objectives is effectively handled by a multi-critic PPO algorithm. In addition, the transformer-based architecture is adaptive to various number of target keyframes and improves accuracy in reaching targets through future goal anticipation.
@@
Acknowledgments
The authors would like to thank Miguel Zamora for valuable discussions. This work was supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (grant agreement No. 866480).
@@ A.1 Observation and Action Space
The observation of the policy is composed of two main components: state observation and goal observation. State observation at time $t$ include the linear velocity ($\bm{v}$) and angular velocity ($\bm{\omega}$) of the base in local coordinates, current joint angles ($\bm{\theta}_{j}$), current joint velocities ($\bm{\dot{\theta}}_{j}$), projected gravity in the base frame ($\bm{g}_{proj}$), base height ($h$) and previous actions from the last timestep ($\bm{a}_{prev})$,
@@ A.2 Reward Terms
Joint soft limits $\mathcal{K}\left(\max\left(\bm{\theta}_{j}-\bm{\theta}_{j,max},\bm{\theta}_{j,min}-\bm{\theta}_{j},\bm{0}\right),0.1,0\right)$
@@
For advantage weights (Equation 2), we recommend starting with a 1:1 ratio and making minor adjustments based on specific preferences. In our experiments, we used $\omega_{style}=0.5,\omega_{goal}=0.5,\omega_{reg}=0.1$, maintaining equal weights for goal and style, with a lower weight for the regularization critic.
@@
Goal position $\Phi^{i}(\mathcal{K}(\bm{p}-\bm{\hat{p}}^{i},0.2,0))$
Goal roll $\Phi^{i}(\mathcal{K}(\phi-\hat{\phi}^{i},0.1,0))$
Goal pitch $\Phi^{i}(\mathcal{K}(\zeta-\hat{\zeta}^{i},0.1,0))$
Goal yaw $\Phi^{i}(\mathcal{K}(\psi-\hat{\psi}^{i},0.3,0))$
Goal posture $\Phi^{i}(\mathcal{K}({\bm{\theta}_{j}-\hat{\bm{\theta}}_{j}^{i}},0.2\sqrt{12},0))$
@@ A.3 Dataset Preparation
We use a database of motion capture from dogs introduced by Zhang et al. [44]. The motions are retargeted to the robot skeleton using inverse kinematics for the end-effectors’ positions with some local offsets to compensate for the different proportions of the robot and dog. A subset of around 20 minutes of data was used, removing the undesired motions such as smelling the ground, walking on slopes, etc. We augment this dataset with other motion clips animated by artists to include more diversity in the dataset. The frame rate is adjusted to that of the simulation, i.e. 50 frames per second. In addition, the linear and angular velocity of the base are precomputed for each frame using a finite difference method and stored in the dataset.
@@ A.4 Training Procedure
We utilize Isaac Gym [55] for simulating the physical environment. At the start of each episode, the robot is initialized either to a default state ($20\%$ probability) or using Reference State Initialization (RSI) ($80\%$ probability). For RSI, a random frame from the reference motion trajectory provides the height, orientation, joint angles, base linear velocity, base angular velocity, and joint velocities for initializing the robot. RSI plays a crucial role in capturing and learning the specific style of motion, as highlighted in previous studies such as Peng et al. [56].
We employ two keyframe sampling strategies: dataset-based and random sampling. In the dataset-based approach, a motion trajectory and starting frame index are initially sampled. Then the time interval between keyframes are uniformly drawn from a specified range. The corresponding frames from the dataset are used to obtain the global position, orientation, and posture of the robot at the keyframe times. The global horizontal positions are subsequently converted to relative local positions with respect to the starting frame and then transformed into positions in the simulator’s world coordinates by accounting for the robot’s initial state. The yaw angle is similarly adjusted, while height, roll, pitch, and joint postures are directly used as target values.
In the random sampling strategy, time intervals are sampled from a predefined range, similar to the first strategy. The radius and direction of each keyframe, relative to the previous one are uniformly sampled to compute the keyframe’s global horizontal position. Additionally, the change in yaw angle is sampled and added to the previous keyframe’s yaw. The roll and pitch angles, joint postures and height are taken directly from a reference frame randomly selected from the dataset. This ensures that the posture is feasible and compatible with the height. Details of the sampling ranges are provided in Table A3.
Our methodology incorporates a learning curriculum, beginning with keyframes entirely sourced from reference data and progressively increasing the proportion of randomly generated keyframes. The meticulous sampling of target keyframes is critical for ensuring their feasibility and preventing them from impeding effective policy learning.
@@
Table A3: Keyframe sampling ranges
Quantity Time interval Waypoint radius Waypoint direction Delta yaw angle
(steps) (m) (rad) (rad)
Range $[25,50]$ $[0.5,1.0]$ $[-\pi/3,\pi/3]$ $[-\pi/3,\pi/3]$
@@ A.5 Hardware Implementation Details
Domain randomization is added during training to achieve a robust policy that can be executed on hardware. Similar to Kang et al. [58], we randomize friction coefficients, motor stiffness and damping gains and actuator latency. Furthermore, we add external pushes during training. Details of domain randomization ranges are provided in Table A4. Although joint limits are softly taken into account in the simulation, we found it crucial to terminate episodes when reaching joint limits to ensure a stable deployment on hardware. We use a motion capture system to receive the global position and orientation of the robot. These are used to compute the relative errors to the target goals and are then passed to the policy. Other observations are computed based on the outputs from the state estimator. Our locomotion control policy runs at $50$ Hz and updates the joint angle targets. At a lower-level, the robot’s built-in motor controller operates at 1000 Hz, updating the torque targets for each joint. To enhance computational efficiency during RL training, we simulate the robot and the PD controller at $200$ Hz, which has proven effective in practice.
Table A4: Domain randomization ranges
Quantity Range
Friction Coefficient $[0.25,1.75]$
Push velocity ($m/s$) $[0.0,1.0]$
Push angular velocity ($m/s$) $[0.0,1.0]$
Stiffness multiplier $[0.9,1.1]$
Damping multiplier $[0.9,1.1]$
Actuator lag steps $[0,6]$
A.6 Keyframe Matching
We provide quantitative analysis of keyframe matching performance for the three scenarios from Figure 5 with full-pose keyframes in Table A5. The mean and standard deviation of each entry is from 20 independent evaluations. The posture error reported is the root mean squared error (RMSE) of all joint angles. The results indicate that our trained policy tracks the keyframes with good precision and consistency.
Table A5: Keyframe matching error for three different scenarios
Scenario Walk (goal 1) Walk (goal 2) Jump Paw up
Distance error (m) $0.048\pm 0.011$ $0.020\pm 0.003$ $0.118\pm 0.012$ $0.044\pm 0.015$
Roll error (deg) $3.094\pm 0.327$ $1.621\pm 0.074$ $1.473\pm 0.579$ $1.209\pm 0.733$
Pitch error (deg) $0.332\pm 0.178$ $0.934\pm 0.235$ $1.558\pm 0.940$ $0.516\pm 0.246$
Yaw error (deg) $1.209\pm 0.384$ $2.120\pm 0.367$ $7.110\pm 1.060$ $2.773\pm 1.301$
Posture error (deg) $2.469\pm 0.309$ $2.595\pm 0.126$ $3.226\pm 0.103$ $4.698\pm 0.808$
To assess the generalization capability of our method, we evaluated the trained policy using three different keyframe sampling strategies:
  • Dataset: Sampling the full pose (position, orientation, and posture) only from trajectories in the reference dataset.
  • Random: Sampling the posture from the reference dataset while randomly sampling the planar position and yaw angle targets.
  • Evaluation dataset: Sampling the full pose from the evaluation dataset, which was not seen by the policy during training. For the evaluation dataset, we use a few locomotion clips from the open-sourced dataset from Han et al. [59].
The mean and standard deviation for each strategy are derived from 20 independent evaluations and shown in Table A6. The results indicate that our policy generalizes outside the trained dataset in terms of position and yaw angle when exposed to the randomly sampled target during training. Surprisingly, the policy achieves decent keyframe tracking in position and orientation as well as postures with the evaluation dataset which is never seen during training. Although Table A6 shows promising generalization results, full generalization to postures that differ significantly from those in the training dataset is yet to be achieved.
Table A6: Keyframe matching performance across different keyframe sampling strategies
Keyframes source Dataset Random Evaluation dataset
Distance error (m) $0.060\pm 0.050$ $0.142\pm 0.087$ $0.064\pm 0.060$
Roll error (deg) $1.873\pm 1.888$ $3.181\pm 3.466$ $3.231\pm 9.053$
Pitch error (deg) $3.084\pm 6.747$ $4.260\pm 9.202$ $1.438\pm 2.842$
Yaw error (deg) $2.394\pm 1.775$ $8.754\pm 6.220$ $2.200\pm 2.235$
Posture error (deg) $5.119\pm 3.143$ $4.640\pm 1.606$ $9.207\pm 7.122$
A.7 Future Goal Anticipation